Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labour-hours, and its standard costs per unit are as follows:

  Direct materials: 5 kg at $8.00 per kg $ 40.00  
  Direct labour: 4 hours at $15 per hour 60.00  
  Variable overhead: 4 hours at $5 per hour 20.00  
  Total standard cost per unit $ 120.00  

The company planned to produce and sell 21,000 units in March. However, during March the company actually produced and sold 24,000 units and incurred the following costs:

  1. Purchased 150,000 kg of raw materials at a cost of $6.40 per kg. All of this material was used in production.
  2. Direct labour: 66,000 hours at a rate of $18 per hour.
  3. Total variable manufacturing overhead for the month was $413,820.

1. What is the materials price variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)

2. What is the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)

5. What is the labour rate variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Do not round intermediate calculations.)

6. What is the labour efficiency variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Do not round intermediate calculations.)

7. What is the variable overhead spending variance for March? (Do not round intermediate calculations. Round the actual overhead rate to two decimal places. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)

8. What is the variable overhead rate variance for March? (Round the actual overhead rate to two decimal places. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)
Answer #1

1) material price variance = (8-6.4)*150000 = 240000 F

2) Material quantity variance = (24000*5-150000)*8 = 240000 U

5) Labor rate variance = (15-18)*66000 = 198000 U

6) Labor efficiency variance = (24000*4-66000)*15 = 450000 F

7) Variable overhead spending variance = (24000*20-413280) = 66720 F

8) Variable overhead rate variance = (5*66000-413280) = 83280 U